### v4

# Hamiltonian stationary Lagrangian surfaces with harmonic mean curvature in complex space forms

Abstract.
In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms whose mean curvature is a harmonic function. We prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel. In the non-constant mean curvature case, under the condition of constant Gaussian curvature, we give an explicit description of all such surfaces.
Key words and phrases:
Hamiltonian stationary Lagrangian surfaces, complex space forms, harmonic mean curvature
2010 Mathematics Subject Classification:
Primary: 53C42; Secondary: 53B25
1. Introduction
[…]
(1.1) ${\rm div}(JH)=0$
on $M$ (cf. [9]), where div is the divergence on $M$.
In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms. First, we prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel (Proposition 3.1). Next, we investigate the case where the mean curvature is a non-constant harmonic function. Our purpose is to give an explicit description of all such surfaces under the condition of constant Gaussian curvature (Theorem 4.1).
2. Preliminaries
Let $\tilde{M}^{n}(4\epsilon)$ be a complete and simply connected complex space form of complex dimension $n$ and constant holomorphic sectional curvature $4\epsilon$, that is, $\tilde{M}^{n}(4\epsilon)$ is the complex Euclidean space $\mathbb{C}^{n}$, the complex projective space $\mathbb{C}P^{n}(4\epsilon)$ or the complex hyperbolic space $\mathbb{C}H^{n}(4\epsilon)$ according as $\epsilon=0$, $\epsilon>0$ or $\epsilon<0$.
Let $M$ be a Lagrangian submanifold of $\tilde{M}^{n}(4\epsilon)$. We denote the Levi-Civita connections on $M^{n}$ and $\tilde{M}^{n}(4\epsilon)$ by $\nabla$ and $\tilde{\nabla}$, respectively. The Gauss and Weingarten formulas are respectively given by
$\tilde{\nabla}_{X}Y=\nabla_{X}Y+h(X,Y),\quad\tilde{\nabla}_{X}\xi=-A_{\xi}X+D_{X}\xi$
for tangent vector fields $X$, $Y$ and normal vector field $\xi$, where $h,A$ and $D$ are the second fundamental form, the shape operator and the normal connection. The mean curvature vector field $H$ is defined by $H=(1/n){\rm trace}\,h.$ The function $|H|$ is called the mean curvature. We have (cf. [7])
(2.1) $D_{X}JY=J(\nabla_{X}Y),$
(2.2) $\left<h(X,Y),JZ\right>=\left<h(Y,Z),JX\right>=\left<h(Z,X),JY\right$>.
Denote by $R$ the Riemann curvature tensor of $\nabla$. Then the equations of Gauss and Codazzi are respectively given by
$\left<R(X,Y)Z,W\right$>= $\left<h(Y,Z),h(X,W)\right>-\left<h(X,Z),h(Y,W)\right$>
[…]
(2.4) $({\bar{\nabla}}_{X}h)(Y,Z)=$ $({\bar{\nabla}}_{Y}h)(X,Z),$
where $X,Y,Z,W$ are vectors tangent to $M$, and $\bar{\nabla}h$ is defined by
$({\bar{\nabla}}_{X}h)(Y,Z)=D_{X}h(Y,Z)-h(\nabla_{X}Y,Z)-h(Y,\nabla_{X}Z).$
3. Hamiltonian stationary Lagrangian surfaces
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H\neq 0$ everywhere. Denote by $K$ the Gaussian curvature of $M$. Let $\{e_{1},e_{2}\}$ be a local orthonormal basis of $M$ such that $Je_{1}$ is parallel to $H$. It follows from (2.2) that second fundamental form takes the form
(3.1) $\begin{split}&h(e_{1},e_{1})=(a-c)Je_{1}+bJe_{2},\\ &h(e_{1},e_{2})=bJe_{1}+cJe_{2},\\ &h(e_{2},e_{2})=cJe_{1}-bJe_{2}\end{split}$
for some functions $a$, $b$ and $c$. Putting $\omega_{i}^{j}(e_{k})=\left<\nabla_{e_{k}}e_{i},e_{j}\right$>, by (2.1) and (LABEL:sf) we have
$(\bar{\nabla}_{e_{1}}h)(e_{2},e_{2})$ $=\{e_{1}c+3b\omega_{1}^{2}(e_{1})\}Je_{1}-\{e_{1}b-3c\omega_{1}^{2}(e_{1})\}Je_{2},$
$(\bar{\nabla}_{e_{2}}h)(e_{1},e_{2})$ $=\{e_{2}b+(a-3c)\omega_{1}^{2}(e_{2})\}Je_{1}+\{e_{2}c+3b\omega_{1}^{2}(e_{2})\}Je_{2},$
$(\bar{\nabla}_{e_{1}}h)(e_{1},e_{2})$ $=\{e_{1}b+(a-3c)\omega_{1}^{2}(e_{1})\}Je_{1}+\{e_{1}c+3b\omega_{1}^{2}(e_{1})\}Je_{2},$
$(\bar{\nabla}_{e_{2}}h)(e_{1},e_{1})$ $=\{e_{2}(a-c)-3b\omega_{1}^{2}(e_{2})\}Je_{1}+\{e_{2}b+(a-3c)\omega_{1}^{2}(e_{2})\}Je_{2}.$
Therefore, the equation (2.4) of Codazzi implies
(3.2) $e_{1}c+3b\omega_{1}^{2}(e_{1})=e_{2}b+(a-3c)\omega^{2}_{1}(e_{2}),$
(3.3) $-e_{1}b+3c\omega_{1}^{2}(e_{1})=e_{2}c+3b\omega^{2}_{1}(e_{2}),$
(3.4) $e_{2}(a-c)-3b\omega_{1}^{2}(e_{2})=e_{1}b+(a-3c)\omega^{2}_{1}(e_{1}).$
Combining (3.3) and (3.4) yields
(3.5) $e_{2}a-a\omega^{2}_{1}(e_{1})=0.$
The Hamiltonian stationary condition (1.1) is equivalent to
(3.6) $e_{1}a+a\omega_{1}^{2}(e_{2})=0.$
Using (3.5) and (3.6), we obtain
$a^{-1}e_{1},a^{-1}e_{2}]=0.$
Therefore, there exists a local coordinate system $\{u,v\}$ such that $e_{1}=a\partial_{u}$, $e_{2}=a\partial_{v}$. Hence, the metric tensor is given by
(3.7) $g=a^{-2}(du^{2}+dv^{2}),$
which implies that
(3.8) $\omega_{1}^{2}(e_{1})=a_{v},\quad\omega_{1}^{2}(e_{2})=-a_{u},$
(3.9) $K=-(a_{u})^{2}-(a_{v})^{2}+a(a_{uu}+a_{vv}).$
Equations (3.2) and (3.3) are rewritten as
(3.10) $ac_{u}+3ba_{v}$ $=ab_{v}-(a-3c)a_{u},$
(3.11) $-ab_{u}+3ca_{v}$ $=ac_{v}-3ba_{u}.$
Put $\delta=\epsilon-K$. Then the equation (2.3) of Gauss and (LABEL:sf) yield that
(3.12) $\delta=2b^{2}-ac+2c^{2}.$
Proposition 3.1.
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$. If $|H|$ is non-zero constant, then $\bar{\nabla}h=0$.
Proof. From (3.5) and (3.6), it follows that if $|H|=|a|/2$ is non-zero constant, then $\omega_{1}^{2}=0$. Hence we have $K=0$. Differentiating (3.12) along $e_{i}$ $(i=1,2)$ yields $4be_{i}b+(4c-a)e_{i}c=0$, which together with (3.2) and (3.3), shows that $b$ and $c$ are constant. Thus, using $\omega_{1}^{2}=0$, (2.1) and (LABEL:sf), we obtain $\bar{\nabla}h=0$.
Remark 3.1.
An explicit description of all Lagrangian surfaces with $\bar{\nabla}h=0$ in $\tilde{M}^{2}(4\epsilon)$ has been obtained in Theorem 7.2 in [1], Theorem A.1 and Theorem A.2 in [4].
4. Main theorem
Let ${\mathbb{C}}^{3}_{1}$ be the complex 33-space endowed with the inner product
$\left<(z_{1},z_{2},z_{3}),(w_{1},w_{2},w_{3})\right>={\rm Re}(-z_{1}\bar{w}_{1}+z_{2}\bar{w}_{2}+z_{3}\bar{w}_{3}).$
Put $H_{1}^{5}(-1)=\{z\in{\mathbb{C}}^{3}_{1}:\left<z,z\right>=-1\}$. The map $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ defined by $z\mapsto z\cdot{\mathbb{C}}^{*}$ is referred to as the Hopf fibration.
The main result of this paper is the following classification theorem.
Theorem 4.1.
Let $M$ be a Hamiltonian stationary Lagrangian surface in a 22-dimensional complex space form $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H$ is nowhere vanishing. If $M$ has non-constant harmonic mean curvature and constant Gaussian curvature $K$, then $K=\epsilon=-1$ and $M$ is locally congruent to the image of $\Pi\circ\phi$, where $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ is the Hopf fibration and $\phi:M\rightarrow H_{1}^{5}(-1)\subset{\mathbb{C}}^{3}_{1}$ is given by one of the following immersions::
$(1)$
$\phi(x,y)=\biggl(me^{y}+\dfrac{e^{-y}+2im^{2}xe^{y}}{2m},me^{ix+y},\dfrac{e^{-y}+2im^{2}xe^{y}}{2m}\biggr);$
$(2)$
$\phi(x,y)=\biggl(1-\dfrac{i(1+m^{2})}{m^{2}x+y},\dfrac{m\sqrt{1+m^{2}}e^{ix}}{m^{2}x+y},\dfrac{\sqrt{1+m^{2}}e^{iy}}{m^{2}x+y}\biggr),$
where $m$ is a positive real number.
Proof. We shall use the same notation as in Section 3. Suppose that $|H|$ is a harmonic function on $M$. Then by (3.7) we have
(4.1) $a_{uu}+a_{vv}=0.$
Moreover, suppose that $|H|$ is non-constant and $K$ is constant. Then, combining (3.9) and (4.1) shows that $K<0$ and
(4.2) $a_{u}=\sqrt{-K}\cos\theta,\quad a_{v}=\sqrt{-K}\sin\theta.$
for some function $\theta$ on $M$. Substituting (4.2) into (4.1), we have
(4.3) $(\sin\theta)\theta_{u}+(\cos\theta)\theta_{v}=0.$
On the other hand, since $a_{uv}-a_{vu}=0$ holds, by (4.2) we obtain
(4.4) $(\cos\theta)\theta_{u}+(\sin\theta)\theta_{v}=0.$
It follows (4.3) from (4.4) that $\theta_{u}=\theta_{v}=0$, that is, $\theta$ is constant. Solving (4.2), we conclude that up to translations, $a$ is given by
(4.5) $a=\sqrt{-K}\{(\cos\theta)u+(\sin\theta)v\}.$
Case (i): $b=0$ on an open subset $\mathcal{U}$. In this case, (3.10) and (3.11) reduce respectively to
(4.6) $(a-3c)a_{u}+ac_{u}=0,$
(4.7) $3ca_{v}-ac_{v}=0.$
Differentiating (3.12) with respect to $v$, we have
(4.8) $ca_{v}+(a-4c)c_{v}=0.$
Combining (4.7) and (4.8) gives
$(a-c)c_{v}=0.$
If $c_{v}\neq 0$ on an open subset in $\mathcal{U}$, then $a=c$ on that open subset. From (3.12) we see that $a$ is constant, which contradicts our assumption. Hence we have $c_{v}=0$ on $\mathcal{U}$. Thus, by (4.7) we get
$ca_{v}=0.$
Case (i.1): $c=0$ on an open subset $\mathcal{U}_{1}\subset\mathcal{U}$. In this case, from (3.12) we have $K=\epsilon=-1$. It follows from (4.6) that $a_{u}=0$ on $\mathcal{U}_{1}$. Hence, using (4.5) and K=1K=-1 yields that $a^{2}=v^{2}$. Applying the coordinate transformation: $y=-\int v^{-1}dv$, we see that the metric tensor (3.7) becomes
$g=m^{2}e^{2y}du^{2}+dy^{2}$
for some positive constant $m$, and the second fundamental form satisfies
$h(\partial_{u},\partial_{u})=J\partial_{u},\quad h(\partial_{u},\partial_{y})=h(\partial_{y},\partial_{y})=0.$
Rewriting $u$ with $x$ and according to [3, page 3475], we conclude that $\mathcal{U}_{1}$ is congruent to the Lagrangian surface obtained from (1).
Case (i.2): $a_{v}=0$ on an open subset $\mathcal{U}_{2}\subset\mathcal{U}$. In this case, $a_{u}\neq 0$. Differentiating (3.12) with respect to $u$ leads to
(4.9) $ca_{u}+(a-4c)c_{u}=0.$
Eliminating $c_{u}$ from (4.6) and (4.9) gives
$(a-2c)(a-6c)a_{u}=0.$
If $a\neq 2c$ on an open set in $\mathcal{U}_{2}$, then $a=6c$ and it follows from (3.12) that $a$ is constant, which contradicts our assumption. Hence $a=2c$, which together with (3.12) implies $K=\epsilon=-1$. Thus, by (4.5) we obtain $a^{2}=u^{2}$. Applying the coordinate transformation: $x=(u+v)/2$ and y=(uv)/2y=(u-v)/2, we see that the metric tensor (3.7) becomes
$g=\dfrac{2}{(x+y)^{2}}(dx^{2}+dy^{2})$
and the second fundamental form satisfies
$h(\partial_{x},\partial_{x})=J\partial_{x},\quad h(\partial_{x},\partial_{y})=0,\quad h(\partial_{y},\partial_{y})=J\partial_{y}.$
According to [5, page 124], $\mathcal{U}_{2}$ is congruent to the Lagrangian surface obtained from (2) with $m=1$.
Case (ii): $b\neq 0$ on an open subset $\mathcal{V}$. We put $A=a_{u}$ and $B=a_{v}$, which are constants. Differentiating (3.12) with respect to $u$ and $v$, we obtain
(4.10) $b_{u}=\frac{cA+(a-4c)c_{u}}{4b},$
(4.11) $b_{v}=\frac{cB+(a-4c)c_{v}}{4b}.$
Substituting (4.10) and (4.11) into (3.10) and (3.11) yields
(4.12) $4abc_{u}-a(a-4c)c_{v}=4b(3c-a)A-(12b^{2}-ac)B,$
(4.13) $a(a-4c)c_{u}+4abc_{v}=(12b^{2}-ac)A+12bcB.$
Solving (4.12) and (4.13) for $c_{u}$ and $c_{v}$ and using (3.12), we have
(4.14) $c_{u}=\frac{Af_{1}+Bf_{2}}{a(a^{2}+8\delta)},$
(4.15) $c_{v}=\frac{Af_{3}+Bf_{4}}{a(a^{2}+8\delta)},$
where
$f_{1}=-4ab^{2}-a^{2}c+4ac^{2},$
$f_{2}=-48b^{3}+16abc-48bc^{2},$
$f_{3}=4a^{2}b+48b^{3}-32abc+48bc^{2},$
$f_{4}=12ab^{2}-a^{2}c+4ac^{2}.$
Substituting (4.14) and (4.15) into (4.10) and (4.11) gives
(4.16) $b_{u}=$ $\frac{Ag_{1}+Bg_{2}}{a(a^{2}+8\delta)},$
(4.17) $b_{v}=$ $\frac{Ag_{3}+Bg_{4}}{a(a^{2}+8\delta)},$
where
$g_{1}=-a^{2}b+8abc,$
$g_{2}=-12ab^{2}+4a^{2}c+48b^{2}c-28ac^{2}+48c^{3},$
$g_{3}=a^{3}+12ab^{2}-12a^{2}c-48b^{2}c+44ac^{2}-48c^{3},$
$g_{4}=3a^{2}b-8abc.$
Differentiating (4.14) and (4.15) with respect to $v$ and $u$ respectively, we have
(4.18) $c_{uv}=\frac{(Af_{1}+Bf_{2})_{v}a(a^{2}+8\delta)-(Af_{1}+Bf_{2})(3a^{2}+8\delta)B}{a^{2}(a^{2}+8\delta)^{2}},$
(4.19) $c_{vu}=\frac{(Af_{3}+Bf_{4})_{u}a(a^{2}+8\delta)-(Af_{3}+Bf_{4})(3a^{2}+8\delta)A}{a^{2}(a^{2}+8\delta)^{2}}.$
By a straightforward but tedious computation, using (4.14)-(4.17) and substituting (3.12) into the numerators of (4.18) and (4.19), we get
(4.20) $c_{uv}=\frac{A^{2}h_{1}+ABh_{2}+B^{2}h_{3}}{a^{2}(a^{2}+8\delta)^{2}},$
(4.21) $c_{vu}=\frac{A^{2}h_{4}+ABh_{5}+B^{2}h_{6}}{a^{2}(a^{2}+8\delta)^{2}},$
where
$h_{1}=$ $-12a^{4}b-144a^{2}b^{3}+160a^{3}bc+768ab^{3}c-656a^{2}bc^{2}+768abc^{3},$
$h_{2}=$ $-108a^{3}b^{2}-960ab^{4}+18a^{4}c+1168a^{2}b^{2}c+2304b^{4}c-252a^{3}c^{2}$
$-3840ab^{2}c^{2}+1296a^{2}c^{3}+4608b^{2}c^{3}-2880ac^{4}+2304c^{5},$
$h_{3}=$ $-96a^{2}b^{3}+768b^{5}-384ab^{3}c+32a^{2}bc^{2}+1536b^{3}c^{2}-384abc^{3}+768bc^{4},$
$h_{4}=$ $-8a^{4}b-96a^{2}b^{3}-768b^{5}+128a^{3}bc+1152ab^{3}c-608a^{2}bc^{2}$
$-1536b^{3}c^{2}+1152abc^{3}-768bc^{4},$
$h_{5}=$ $-92a^{3}b^{2}-192ab^{4}+18a^{4}c+784a^{2}b^{2}c+2304b^{4}c-252a^{3}c^{2}$
$-3072ab^{2}c^{2}+1296a^{2}c^{3}+4608b^{2}c^{3}-2880ac^{4}+2304c^{5},$
$h_{6}=$ $-240a^{2}b^{3}+80a^{3}bc+768ab^{3}c-496a^{2}bc^{2}+768abc^{3}.$
Since $c_{uv}-c_{vu}$=0 holds, from (4.20), (4.21), it follows that
(4.22) $4b(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})=0,$
where
$k_{1}=$ $-a^{4}-12a^{2}b^{2}+192b^{4}+8a^{3}c-96ab^{2}c-12a^{2}c^{2}$
$+384b^{2}c^{2}-96ac^{3}+192c^{4},$
$k_{2}=$ $-4a^{3}b-192ab^{3}+96a^{2}bc-192abc^{2},$
$k_{3}=$ $36a^{2}b^{2}+192b^{4}-20a^{3}c-288ab^{2}c$
$+132a^{2}c^{2}+384b^{2}c^{2}-288ac^{3}+192c^{4}.$
By a straightforward but tedious computation, using (4.14)-(4.17), we get
(4.23) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{u}$ $=\frac{4(A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4})}{a^{2}+8\delta},$
(4.24) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{v}$ $=\frac{8(A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8})}{a^{2}+8\delta},$
where
$P_{1}=$ $-a^{5}-24a^{3}b^{2}-192ab^{4}+12a^{4}c+168a^{2}b^{2}c+384b^{4}c$
$-56a^{3}c^{2}-576ab^{2}c^{2}+168a^{2}c^{3}+768b^{2}c^{3}-384ac^{4}+384c^{5},$
$P_{2}=$ $-2a^{4}b-72a^{2}b^{3}-1920b^{5}+24a^{3}bc+1344ab^{3}c$
$-264a^{2}bc^{2}-3840b^{3}c^{2}+1344abc^{3}-1920bc^{4},$
$P_{3}=$ $32a^{3}b^{2}+960ab^{4}-14a^{4}c-888a^{2}b^{2}c-2688b^{4}c+224a^{3}c^{2}$
$+4032ab^{2}c^{2}-1272a^{2}c^{3}-5376b^{2}c^{3}+3072ac^{4}-2688c^{5},$
$P_{4}=$ $24a^{2}b^{3}+1152b^{5}-8a^{3}bc-960ab^{3}c$
$+216a^{2}bc^{2}+2304b^{3}c^{2}-960abc^{3}+1152bc^{4},$
$P_{5}=$ $a^{4}b+60a^{2}b^{3}+576b^{5}-32a^{3}bc-672ab^{3}c$
$+252a^{2}bc^{2}+1152b^{3}c^{2}-672abc^{3}+576bc^{4},$
$P_{6}=$ $-a^{5}-38a^{3}b^{2}-192ab^{4}+24a^{4}c+444a^{2}b^{2}c+1344b^{4}c-218a^{3}c^{2}$
$-2016ab^{2}c^{2}+924a^{2}c^{3}+2688b^{2}c^{3}-1824ac^{4}+1344c^{5},$
$P_{7}=$ $-4a^{4}b-180a^{2}b^{3}-960b^{5}+96a^{3}bc+1248ab^{3}c$
$-564a^{2}bc^{2}-1920b^{3}c^{2}+1248abc^{3}-960bc^{4},$
$P_{8}=$ $6a^{3}b^{2}-5a^{4}c-84a^{2}b^{2}c-192b^{4}c+50a^{3}c^{2}$
$+288ab^{2}c^{2}-180a^{2}c^{3}-384b^{2}c^{3}+288ac^{4}-192c^{5}.$
Thus, it follows from (4.22), (4.23) and (4.24) that
(4.25) $A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4}=0,$
(4.26) $A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8}=0.$
Using a computer algebra system, we check that the resultant of the left-hand sides of (4.25) and (4.26) with respect to AA is, up to a constant factor,
(4.27) $B^{9}(2b^{2}-ac+2c^{2})^{3}(a^{2}+16b^{2}-8ac+16c^{2})^{7}(a^{2}+48b^{2}-24ac+48c^{2})^{2}Q,$
where
$Q=$ $5a^{6}+120a^{4}b^{2}+720a^{2}b^{4}+2304b^{6}-60a^{5}c-720a^{3}b^{2}c$
$-3456ab^{4}c+300a^{4}c^{2}+3168a^{2}b^{2}c^{2}+6912b^{4}c^{2}-1008a^{3}c^{3}$
$-6912ab^{2}c^{3}+2448a^{2}c^{4}+6912b^{2}c^{4}-3456ac^{5}+2304c^{6}.$
Combining (3.12) and (4.27), we find that (4.27) can be simplified to
(4.28) $B^{9}\delta^{3}(a^{2}+8\delta)^{7}(a^{2}+24\delta)^{2}(5a^{6}+60a^{4}\delta+180a^{2}\delta^{2}+288\delta^{3}),$
which must vanish on $\mathcal{V}$. Taking into account $\nabla a\neq 0$, we have
$B\delta=0.$
If $\delta\neq 0$, then $B=0$, which implies that (4.22) and (4.26) reduce respectively to
$a^{4}+12a^{2}b^{2}-192b^{4}-8a^{3}c+96ab^{2}c+12a^{2}c^{2}$
(4.29) $-384b^{2}c^{2}+96ac^{3}-192c^{4}=0,$
$a^{4}+60a^{2}b^{2}+576b^{4}-32a^{3}c-672ab^{2}c$
(4.30) $+252a^{2}c^{2}+1152b^{2}c^{2}-672ac^{3}+576c^{4}=0.$
Subtracting (4.29) from (4.30), we derive
(4.31) $24(2b^{2}-ac+2c^{2})(a^{2}+16b^{2}-8ac+16c^{2})=0.$
Combining (3.12) and (4.31) implies
$a^{2}+8\delta=0,$
which contradicts $\nabla a\neq 0$. Hence $\delta=0$, that is, $K=\epsilon=-1$.
From (3.12) we see that $c\neq 0$. Changing the sign of $e_{1}$ if necessary, we may assume that $c>0$. Put
(4.32) $f=\frac{\sqrt{b^{2}+c^{2}-b\sqrt{b^{2}+c^{2}}}}{\sqrt{2}(b^{2}+c^{2})},\quad k=\frac{\sqrt{b^{2}+c^{2}+b\sqrt{b^{2}+c^{2}}}}{\sqrt{2}(b^{2}+c^{2})},$
which are non-zero and unequal everywhere. It follows from (3.12) and (4.32) that
(4.33) $a=\frac{\sqrt{f^{2}+k^{2}}}{fk},\quad b=\frac{k^{2}-f^{2}}{(f^{2}+k^{2})^{\frac{3}{2}}},\quad c=\frac{2fk}{(f^{2}+k^{2})^{\frac{3}{2}}}.$
We make the following change of the basis:
(4.34) $\tilde{e}_{1}=\frac{k}{\sqrt{f^{2}+k^{2}}}e_{1}+\frac{f}{\sqrt{f^{2}+k^{2}}}e_{2},\quad\tilde{e}_{2}=\frac{f}{\sqrt{f^{2}+k^{2}}}e_{1}-\frac{k}{\sqrt{f^{2}+k^{2}}}e_{2}.$
Then, using (LABEL:sf), (4.33) and (4.34), we find
$h(\tilde{e}_{1},\tilde{e}_{1})=f^{-1}J\tilde{e}_{1},\quad h(\tilde{e}_{1},\tilde{e}_{2})=0,\quad h(\tilde{e}_{2},\tilde{e}_{2})=k^{-1}J\tilde{e}_{2}.$
Therefore, according to [5, Theorem 4.1], there exists a local coordinate system $\{x,y\}$ such that $\partial_{x}=f\tilde{e}_{1}$ and $\partial_{y}=k\tilde{e}_{2}$, where $f$ and $k$ satisfy
(4.35) $\frac{f_{y}}{k}=\frac{k_{x}}{f},\quad\Bigl(\frac{f_{y}}{k}\Bigr)_{y}+\Bigl(\frac{k_{x}}{f}\Bigr)_{x}=fk.$
The Hamiltonian stationary condition (1.1) is equivalent to (cf. [8])
(4.36) $\Bigl(\frac{k}{f}\Bigr)_{x}+\Bigl(\frac{f}{k}\Bigr)_{y}=0.$
By Theorem 3.1 of [2], up to translations and sign, the exact solutions of the over-determined PDE system (4.35)-(4.36) are given by
(4.37) $f=\lambda m\,{\rm csch}\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr),\quad$ $k=\lambda\,{\rm csch}\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr);$
(4.38) $f=\lambda m\,\sec\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr),\quad$ $k=\lambda\,\sec\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr);$
(4.39) $f=\frac{m\sqrt{1+m^{2}}}{m^{2}x+y},\quad$ $k=\frac{\sqrt{1+m^{2}}}{m^{2}x+y},$
where $\lambda$ and $m$ are positive real numbers.
We note that $f=mk$. By the first equation of (4.33) we have
$a=\frac{\sqrt{1+m^{2}}}{mk}.$
Furthermore, it follows from (4.34) that
$\partial_{u}=\frac{1}{1+m^{2}}(\partial_{x}+m^{2}\partial_{y}),\quad\partial_{v}=\frac{m}{1+m^{2}}(\partial_{x}-\partial_{y}).$
Since $a_{u}$ and $a_{v}$ are constant, the solutions (4.37) and (4.38) are excluded. Considering also that $a_{u}^{2}+a_{v}^{2}=1$ and $f\neq k$, we see that $f$ and $g$ are given by (4.39) with $m\neq 1$. In [6, Section 7], it is proved that the corresponding surface is congruent to the Lagrangian surface obtained from case (2) with $m\neq 1$. This completes the proof.